OpenStudy (anonymous):

solve. 0.5x+3(is greater than or equal to)1.3x-6 the solution is {xIx__ __

OpenStudy (saifoo.khan):

\[0.5x + 3 \ge 1.3x -6\]

OpenStudy (anonymous):

i do add 100 to all correct?

OpenStudy (saifoo.khan):

\[1.3x -6 \le 0.5x + 3\]\[0.8x \le 9\]\[x \le \frac{9}{0.8}\]\[x \le 11.25\]

OpenStudy (anonymous):

Im just lost in this class.

OpenStudy (saifoo.khan):

\[x|x \le 11.25\]

OpenStudy (saifoo.khan):

are u doing online school? I can help where are u stuck?

OpenStudy (anonymous):

im doing online college. My whole problem is retention.

OpenStudy (anonymous):

Its my final that im working on and if I dont pass the final I fail

OpenStudy (saifoo.khan):

i have done this like 3-4 times.

OpenStudy (saifoo.khan):

come here, i will help you out and show u step by step. =)